# Infinitely many hyperelliptic curves of small genus and small fixed rank, and of any genus and rank two

### v3

@@ 3.1. Higher genus - over any number field
The easier case is when g=2k1g=2k-1. We may assume that $a_{2k+1}=0$ (otherwise just replace $x-a_{2k+1}$ by $x$). We need to glue the curve $C$ with $Y:y^{2}=(x-a_{1})\cdots(x-a_{2k})$. We choose $B>0$ from [KM, Theorem 1.3] for genus k1k-1. Our assumption of being BB-über-generic clearly implies that $Y$ is $B$-generic. Hence, by [KM, Theorem 1.3], there are infinitely many twists $Y_{d}:dy^{2}=(x-a_{1})\cdots(x-a_{2k})$ of $Y$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))=1$. Then, for $C_{d}:dy^{2}=(x^{2}-da_{1})\cdots(x^{2}-da_{2k})$, by Lemma 3.1, we have that $\mathrm{Jac}(C_{d})$ is isogenous to $\mathrm{Jac}(X)\times\mathrm{Jac}(Y_{d})$ over $K$. Hence, $\mathrm{Rank}(\mathrm{Jac}(C_{d})(K))=r+1$, and the argument that there are infinitely many non-isomorphic curves among them is as before.
Now, we consider the case $g=2k$. We may assume now $a_{1}\cdots a_{2k+1}\neq 0$. We need to glue the curve $C$ with $Y:y^{2}=x(x-a_{1})\cdots(x-a_{2k+1})$. We choose $B>0$ from [KM, Theorem 1.3] for genus $k$. We need an odd model of $Y$, so the change of variables $x\mapsto\frac{1}{x}$, induces that $Y$ is isomorphic over $K$ to
@@
so $Z$ is $B$-generic. Therefore, we can apply [KM, Theorem 1.3] to conclude that there are infinitely many twists $Z_{d}$ of $Z$ such that $\mathrm{Rank}(\mathrm{Jac}(Z_{d})(K))=1=\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))$ for the corresponding twist $Y_{d}$ of $Y$. Now, we conclude as in the previous case. ∎
@@
Let $g\geq 3$ be an integer and $K$ a number field. There are infinitely many hyperelliptic curves $C$ such that $\mathrm{Rank}(\mathrm{Jac}(C)(K))=r$ for $r=1$ and $2$.
@@
By [KM], we can always find a quadratic twist of a BB-über-generic curve $X$ of genus $\lfloor\frac{g+1}{2}\rfloor$ whose Jacobian over $K$ has rank $1$. We then apply Theorem 3.10 (++ Remark 3.6) and Theorem 3.13 to conclude. ∎
@@ 3.2. Examples over $\mathbb{Q}$
Now we prove the existence of infinitely many curves over $\mathbb{Q}$ of genus $g$ whose Jacobian over $\mathbb{Q}$ as rank $r$ for $g\in\{3,4\}$ and $1\leq r\leq 4$ and $g\in\{5,6\}$ and $1\leq r\leq 3$. Even though the existence of the rank 1 and 2 follows from Theorem 3.15, we still decided to include these examples.
@@ 3.2.2. Genus five and six
Even though it sounds appealing to use Theorem 3.13 to construct examples, we note that the $B$-generic and BB-über-generic conditions are less practical to use. Namely, a careful inspection on the condition on $B$ from [KM] shows that $B>9600$ for genus two curves, so we will need to use several large primes to construct such $B$-generic or BB-übergeneric curves. Moreover, solving the Chinese Reminder Theorem system will result in large coefficients of the curve $C$. Hence, we have not tried to use $B$-generic or BB-übergeneric curves to get concrete examples.
@@ 4. An answer to our question
If the Jacobian of a curve is decomposable as a product of Jacobians of curves, then we reduce the problem of computing the rank to smaller genus curves, which is, in principle, way easier. The problem becomes much harder if we restrict the question for curves $C$ of any genus such that $\mathrm{Jac}(C)$ is absolutely simple. The groundbreaking work of Peter Koymans and Adam Morgan [KM] confirms that there are infinitely many absolutely simple Jacobians of hyperelliptic curves such that the rank of $\mathrm{Jac}(C)(\mathbb{Q})$ is equal to $1$. This result resolves the following question, which was posed as a challenge in the previous version of this article as stated in June 2025.
@@
We keep the question as stated in its original form, as [KM, Theorem 1.1] specifically cites the problem mentioned beforehand. We sincerely congratulate and thank them for solving the question stated above.
@@ Acknowledgements
We would like to thank the Max Planck Institute for Mathematics for its generous support. We would like to also thank Lodha Mathematical Sciences Institute for its generous support and hospitality during the Thematic Programme on Arithmetic Statistics where substantial results of this paper were achieved. We would especially like to thank Jennifer Balakrishnan, Maarten Derickx, Dušan Dragutinović, Peter Koymans, Adam Morgan, Carlo Pagano, and Lazar Radičević for numerous conversations and their enlightening comments about this work. We would like also to thank Jordan Ellenberg, Matija Kazalicki, Jef Laga, Filip Najman, Oana Padurariu, Alex Smith, and Wadim Zudilin for helpful discussions.

### v4

@@ 3.3. Our constructions
The easier case is when g=2k1g=2k-1. We may assume that $a_{2k+1}=0$ (otherwise just replace $x-a_{2k+1}$ by $x$). We aim to create the curve $C$ as a fibre product of $X_{m}$ with $Y_{dm}:{dm}y^{2}=(x-a_{1})\cdots(x-a_{2k})$ (we are in the situation of Equation (4), but with changed twisting numbers, here $m$ corresponds to $dm$ from Equation (4), and then $dm$ here corresponds to $d^{2}m$ equal to $m$ modulo squares from Equation (4)). We choose $B>0$ from [23, Theorem 1.3] for genus k1k-1. Our assumption of being BB-über-generic clearly implies that $Y:y^{2}=(x-a_{1})\cdots(x-a_{2k})$ is $B$-generic. Hence, by [23, Theorem 1.3], there are infinitely many twists $Y_{dm}:{dm}y^{2}=(x-a_{1})\cdots(x-a_{2k})$ of $Y$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{dm})(K))=1$. Then, for $C:my^{2}=(x^{2}-da_{1})\cdots(x^{2}-da_{2k})$, by Remark 3.2, we have that $\mathrm{Jac}(C)$ is isogenous to $\mathrm{Jac}(X_{m})\times\mathrm{Jac}(Y_{dm})$ over $K$. Hence, $\mathrm{Rank}(\mathrm{Jac}(C)(K))=r+1$, and the argument that there are infinitely many non-isomorphic curves among them is as before.
Now, we consider the case $g=2k$. We may assume now $a_{1}\cdots a_{2k+1}\neq 0$. We aim to construct the curve $C$ as a fibre product of $X_{m}$ with $Y_{dm}:dmy^{2}=x(x-a_{1})\cdots(x-a_{2k+1})$ (we are in the situation of Equation (5), but, again, with changed twisting numbers, here $m$ corresponds to $dm$ from Equation (5), and then $dm$ here corresponds to $d^{2}m$ equal to $m$ modulo squares from Equation (5)). We choose $B>0$ from [23, Theorem 1.3] for genus $k$. Now, we prove that $Y:y^{2}=x(x-a_{1})\cdots(x-a_{2k+1})$ is BB-über-generic, so we need to use an odd-degree model for $Y$. The change of variables $x\mapsto\frac{1}{x}$, induces that $Y$ is isomorphic over $K$ to
@@
so $Z$ is $B$-generic. Therefore, we can apply [23, Theorem 1.3] to conclude that there are infinitely many twists $Z_{dm}$ of $Z$ such that $\mathrm{Rank}(\mathrm{Jac}(Z_{dm})(K))=1=\mathrm{Rank}(\mathrm{Jac}(Y_{dm})(K))$ for the corresponding twist $Y_{dm}$ of $Y$. Now, we conclude as in the previous case. ∎
@@
Let $g\geq 3$ be an integer and $K$ a number field. There are infinitely many hyperelliptic curves $C$ of genus $g$ such that $\mathrm{Rank}(\mathrm{Jac}(C)(K))=r$ for $r=1$ and $2$.
@@
By [23], we can always find a quadratic twist of a BB-über-generic curve $X$ of genus $\lfloor\frac{g+1}{2}\rfloor$ whose Jacobian over $K$ has rank $1$. We then apply Theorem 3.10 (++ Remark 3.6) and Theorem 3.13 to conclude. ∎
@@ 3.4. Examples over $\mathbb{Q}$
Now we prove the existence of infinitely many curves over $\mathbb{Q}$ of genus $g$ whose Jacobian over $\mathbb{Q}$ has rank $r$ for $g\in\{3,4\}$ and $1\leq r\leq 4$ and $g\in\{5,6\}$ and $1\leq r\leq 3$. Even though the existence of the rank 1 and 2 follows from Theorem 3.15, we still decided to include these examples.
@@ 3.4.2. Genus five and six
Even though it sounds appealing to use Theorem 3.13 to construct examples, we note that the $B$-generic and BB-über-generic conditions are less practical to use. Namely, a careful inspection on the condition on $B$ from [23] shows that $B>9600$ for genus two curves, so we will need to use many large primes to construct such $B$-generic or BB-über-generic curves. Moreover, solving the Chinese Remainder Theorem system will result in large coefficients of the curve $C$. Hence, we have not tried to use $B$-generic or BB-über-generic curves to get concrete examples.
@@ 4. An answer to our question
If the Jacobian of a curve is decomposable as a product of Jacobians of curves, then we reduce the problem of computing the rank to smaller genus curves, which is, in principle, easier, even though, as we have seen, it also contains its own subtleties. Since we gave a construction of a decomposable Jacobian of a fixed positive rank, we posed a challenge in the previous version of this article as stated in June 2025 to find such examples of curves $C$ such that $\mathrm{Jac}(C)$ is absolutely simple. The breakthrough work of Koymans and Morgan [23] resolves the following question (and in fact for any genus $g$).
@@
We keep the question as stated in its original form, as [23, Theorem 1.1] specifically cites the problem mentioned beforehand. We sincerely congratulate and thank them for solving the question stated above.
@@ Acknowledgements
We would like to thank the Max Planck Institute for Mathematics for its generous support. We would like to also thank Lodha Mathematical Sciences Institute for its generous support and hospitality during the Thematic Programme on Arithmetic Statistics where substantial results of this paper were achieved. We would especially like to thank Maarten Derickx and Carlo Pagano for their advice and inspiring discussions about this project, and in particular, about motivating us to consider the general situation from §3.1, Jennifer Balakrishnan, Dušan Dragutinović, Peter Koymans, Adam Morgan, and Lazar Radičević for numerous conversations and their enlightening comments about this work. We would also like to thank Jordan Ellenberg, Matija Kazalicki, Jef Laga, Filip Najman, Oana Padurariu, Alex Smith, and Wadim Zudilin for helpful discussions.